OpenStudy (praxer):

Find all$$f:R→R$$such that$$ f(x^2+f(y))=(f(x))^2+y,∀ x,y∈R$$ Help me to understand this type of questions.I am new to this pattern. Thanks in advance!

OpenStudy (praxer):

@rational

rvc (rvc):

@Michele_Laino @rational please help :)

OpenStudy (amriju):

is the answer (x+4)/3?

OpenStudy (amriju):

hello..

OpenStudy (praxer):

answer is$$ f(x)=x$$

OpenStudy (amriju):

ok..

rvc (rvc):

@Michele_Laino hey will you please help

OpenStudy (michele_laino):

here I'm

rvc (rvc):

yes will you please wait for sometime Michele please :)

OpenStudy (michele_laino):

ok!

OpenStudy (praxer):

okay I am here :) Thanks sister :) @rvc

rvc (rvc):

Thanks @Michele_Laino for waiting so patiently :)

OpenStudy (praxer):

@Michele_Laino will the function be an indentity function.

OpenStudy (michele_laino):

I think that your equation can be viewed as an operational equation

OpenStudy (michele_laino):

namely a functional equation

OpenStudy (praxer):

yes it is from the olympiad prep materials. And a question from the functional equation section.

OpenStudy (michele_laino):

there exist an operator, say T, such that, your equation can be written as below: \[T\left( f \right) = ...\] so, we have to know who is our operator T

OpenStudy (praxer):

taking x=0 we have $$\begin{align} \forall x,y\in \Bbb R^2&& f\left(f\left(y\right)\right)=y+f(0)^2=y\tag2\end{align}$$ which means that f is one one and onto. can it be ???

OpenStudy (michele_laino):

I'm not sure

OpenStudy (praxer):

for y=o we will have$$ f(x^2)=(f(x))^2$$

OpenStudy (michele_laino):

why f(0)=0?

OpenStudy (michele_laino):

please wait a moment, someone is calling me to my phone

OpenStudy (michele_laino):

here I'm

OpenStudy (michele_laino):

using the Taylor expansion, at the left side, we can write: \[\Large f\left( {{x^2} + f\left( y \right)} \right) = f\left( {{x^2}} \right) + f\left( y \right)\frac{{df}}{{dx}}\]

OpenStudy (michele_laino):

sorry, that formula is wrong

OpenStudy (praxer):

I guess so ! I am confused with the step.

OpenStudy (michele_laino):

for example one solution of your problem is the identity function, as you stated before.